# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ 3.3 Computational Complexity
What is the computational complexity of implementing the rules $\textsc{Cons}_{{\sf e2e}}$ and $\textsc{Cons}_{{\sf CoT}}$, and thus of enjoying the learning guarantees in the preceding theorems? Recall from Section 2 that in order to talk about computational tractability, we consider a hierarchy of base classes ${\mathcal{F}}_{d}$ with some size parameter $d$. If the base class ${\mathcal{F}}_{d}$ itself is computationally hard to learn, we cannot generally hope for the iterated classes to be easier. But what can we say if ${\mathcal{F}}_{d}$ itself is computationally easy to learn? Can we then implement $\textsc{Cons}_{{\sf e2e}}$ and $\textsc{Cons}_{{\sf CoT}}$ efficiently?
@@
Consider any hypothesis class ${\mathcal{F}}$ over an alphabet set $\Sigma$. The rule $\textsc{Cons}_{{\sf CoT}}$ on $m$ examples of input length at most $n$ can be implemented by a single call to $\textsc{Cons}_{\mathcal{F}}$, with $\tilde{m}\leq m\cdot T$ samples of input length at most $n+T$, and $O(\tilde{m})$ additional runtime.
@@
For a family ${\mathcal{F}}_{d}$, if $\textsc{Cons}_{{\mathcal{F}}_{d}}$ can be implemented in time polynomial in its input and the size parameter $d$, and $\mathrm{VCdim}({\mathcal{F}}_{d})\leq{\mathcal{P}oly}(d)$, then ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf CoT}$ learnable in time ${\mathcal{P}oly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$.
In the next section, we will see that the situation is much grimmer for ${\sf e2e}$-learnability. Even for tractably learnable base classes, where $\textsc{Cons}_{\mathcal{F}}$ is computationally easy, implementing $\textsc{Cons}_{{\sf e2e}}$, or in fact ${\sf e2e}$-learning using any other possible rule, could be computationally hard. This computational gap is perhaps the biggest advantage of Chain-of-Thought training. In Sections 5 and 6, we will further see how to leverage the computational tractability of ${\sf CoT}$ learning exposed by Corollary 3.7.
@@ 4 Autoregressive Linear Thresholds
From Theorems 3.2 and 3.4, we have that $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf e2e}$ and ${\sf CoT}$ learnable with sample complexities $m^{\sf e2e}\propto O(Td)$ and $m^{\sf CoT}\propto O(d\log T)$ respectively, with the disappointing $O(T)$ scaling for ${\sf e2e}$ learning. But in this case, even though $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}$ is specified in terms of real-valued parameters, the discreteness of the input domain actually allows us to bound its cardinality:
@@
Plugging in Lemma 4.1 into Theorem 3.1 we see that ${\sf e2e}$ learning is also possible with $T$-independent sample complexity. Combining the VC-dimension and cardinality-based sample complexities:
@@
While we do not know if the $d^{2}$ vs. $d\log T$ gap in the sample complexity (when $\log T\ll d$) between ${\sf e2e}$ and ${\sf CoT}$ learning is real, it is clear that there remains a statistical advantage over learning the time-dependent $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ (as in Malach, 2023), which requires $\Omega(d\cdot T)$ samples. The main advantage, however, of ${\sf CoT}$ learning over ${\sf e2e}$ learning here is computational, rather than statistical. Since the consistency problem $\textsc{Cons}_{\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}}$ amounts to Linear Programming feasibility and is thus easily solvable in polynomial time (Dantzig, 2016), Corollary 3.7 ensures that we can ${\sf CoT}$ learn $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ with $O(d\log T\wedge d^{2})$ samples and implement the rule on inputs of length $n$ in time ${\mathcal{P}oly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log(\delta^{\scriptscriptstyle{-1}}))$:
@@
The class $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf CoT}$-learnable in time ${\mathcal{P}oly}(d,T,n,1/\varepsilon,\log(1/\delta))$.
@@
Computational Complexity Not learnable in ${\mathcal{P}oly}(d,T,n)$, Learnable in ${\mathcal{P}oly}(d,T,n)$,
@@
Table 2: A summary of the results in Section 4 about learnability of time-invariant iterated linear thresholds ${\mathcal{L}}^{{\sf e2e}\textnormal{-}{T}}$. Sample complexities follow Corollary 4.2. We do not know whether these bounds are tight, and it remains open whether the gap between $m^{{\sf e2e}}$ and $m^{{\sf CoT}}$ is real. In Corollary 4.3 and Theorem 4.4, we show the tractability of ${\sf CoT}$-learning and hardness of ${\sf e2e}$-learning of $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ respectively.
On the other hand, we show that $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is not tractably ${\sf e2e}$-learnable, i.e. when Chain-of-Thought is not provided. Not only is $\textsc{Cons}_{{\sf e2e}}$ hard to implement, but we prove that no learning rule can ${\sf e2e}$-learn $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ in polynomial time (even improperly). To show this, we rely on the hardness of learning constant depth threshold circuits—a well-established hardness assumption in learning theory. See Section C.1 for a formal definition of the class of linear threshold circuits.
@@
There exists a constant $L>0$ and a polynomial $p(n)$ such that threshold circuits of depth $L$ and size $p(n)$ over $n$ binary inputs are not weakly PAC-learnable in time ${\mathcal{P}oly}(n)$, i.e. to constant accuracy and confidence parameters $\varepsilon<0.5,\delta<1$.
Assumption 1 is implied by core cryptographic assumptions such as hardness of the problems of inverting the RSA encryption function, recognizing quadratic residues, and factoring Blum integers (see Kearns and Valiant, 1994, Theorem 6). The Decisional Diffie-Hellman (DDH) assumption implies hardness already of learning depth-4 circuits, and so also implies Assumption 1 (Naor and Reingold, 1997, Krause and Lucks, 2001). An intersection of halfspaces is simply a depth-two linear threshold circuit, and hence hardness of learning the intersection of polynomially many halfspaces also implies Assumption 1. This is in turn implied by hardness of the unique shortest vector problem (Klivans and Sherstov, 2009), hardness of refuting random $K$-SAT formulas (Daniely and Shalev-Shwartz, 2016), and the existence of local pseudo-random generators (Daniely and Vardi, 2021).
@@
Under Assumption 1, there is no algorithm that ${\sf e2e}$-learns ${\mathcal{F}}_{\scriptscriptstyle{d,{\rm lin}}}^{{\sf e2e}\textnormal{-}{T}}$ over $\{0,1\}^{n}$ in time ${\mathcal{P}oly}(d,T,n)$ (even to within constants $\varepsilon<0.5,\delta<1$).
Theorem 4.4 follows from a reduction from circuits to time-invariant iterated linear thresholds: we show that any linear threshold circuit with depth $L$ and size $s$ can be emulated using a time-invariant autoregressive linear threshold of dimension $d=O(s^{L})$, up to a fixed poly-time feature map:
@@
In Lemma 4.5, the insistence on time-invariance, i.e. using the same linear threshold function in every step of generation, makes the construction much trickier. If instead we allowed a different linear predictor at each step of generation, as in $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ defined in Eqs. (4)–(5), we could directly encode each node in the circuit as a step of generation, with $T=s$ generation steps and dimensionality $d=s+n$. This was the approach taken by Malach (2023) when studying time-dependent autoregressive generation. We leave it as an open question whether it is possible to improve the time-invariant construction of Lemma 4.5, and reduce the required dimensionality, and especially the dependence on depth. In any case, Lemma 4.5 is sufficient for obtaining the hardness result of Theorem 4.4.999Improving the construction will only enable basing hardness on slightly weaker assumptions, such as hardness of learning log-depth circuits, or even existence of one-way functions.
@@ 5 Expressivity and Universality
One might possibly view Lemma 4.5 as a positive result about the expressive power of autoregressive time-invariant linear thresholds, suggesting they are “universal” and can express any computable function. In this sense Lemma 4.5 is disappointing, yielding a sample complexity of $\text{size}^{\Omega(\text{depth})}$ for learning circuits. Not only is this exponential in depth, but even for very shallow circuits, e.g. $L=2$, the sample complexity is much larger than the size of the circuit and of the sample complexity of learning the circuit directly.
@@
Should this motivate us to attempt and improve the construction in Lemma 4.5? If our goal is to express circuits, the time-dependent class $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ is more appropriate, and we will not be able to improve over the sample complexity it provides for learning circuits (Malach, 2023, Bartlett and Maass, 2003). Circuits are inherently not “time-invariant”, and their size and number of parameters are always (at least) linear in the input length, and more importantly, in the time of the computation involved.
What form of “universal” expressive power do we want then? What would we want the sample complexity to depend on?
@@
One type of desirable universality is to learn all computable functions with sample complexity proportional to the runtime of computing them. Said differently, consider the class $\mathsf{TIME}(\mathtt{T})$ of functions computable in time $\mathtt{T}$. Can we learn $\mathsf{TIME}(\mathtt{T})$ using ${\mathcal{P}oly}(\mathtt{T})$ samples? All such functions can be expressed as circuits of size $\tilde{O}(\mathtt{T})$ (Arora and Barak, 2009), and hence feed-forward neural nets with $E={\mathcal{P}oly}(\mathtt{T})$ edges. Since, from a sample complexity perspective, neural nets (e.g. with threshold activations) can be learned from $\tilde{O}(E)$ samples (Anthony and Bartlett, 2009), neural nets already provide for “universal” learning in this sense: minimizing the loss on a neural net of size ${\mathcal{P}oly}(\mathtt{T})$ would allow us to learn $\mathsf{TIME}(\mathtt{T})$ using ${\mathcal{P}oly}(\mathtt{T})$ samples. The important caveat here, of course, is that learning a neural network is not computationally tractable, and we cannot actually minimize the loss on a neural net in polynomial time.
@@
Time-Dependent Autoregressive Chain-of-Thought learning with linear thresholds, i.e. using $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ is also “universal” in the same way as feed-forward networks: it allows expressing the $\mathsf{TIME}(\mathtt{T})$ with dimensionality and generation length $d,T=O(\mathtt{T})$, and thus learning with ${\mathcal{P}oly}(\mathtt{T})$ samples. The advantage here, as pointed out by Malach (2023), is that we can consider Chain-of-Thought training as essentially providing us a peek into the inner working of the computation, and thus allowing computationally tractable learning.
@@
But in principle, we can hope for much stronger universality than just learning with number of samples proportional to computation runtime. Consider the much broader class ${\sf PROG}(\mathtt{S})$ of programs of length at most $\mathtt{S}$. Many very-long-runtime functions have a short program but require a very long time to run, with runtime that usually depends on input length while program length does not. This is also a “time invariant” class, since the program is independent of the input and computation length. The key to the universality dream in machine learning is that this amazingly universal class has log-cardinality of only $\log{\left\lvert{{\sf PROG}(\mathtt{S})}\right\rvert}=O(\mathtt{S})$ and thus could, in principle, be learned with $O(\mathtt{S})$ samples. The caveat again is intractability of training. We really have no hope of learning ${\sf PROG}(\mathtt{S})$ tractably–not only is finding a short program with small error uncomputable, but also functions in ${\sf PROG}(\mathtt{S})$ might require unbounded amount of time to even evaluate, let alone learn. Such functions are anyway useless as predictors, and it would be sensible to exclude them from consideration in the first place.
@@
This leads us to consider the class ${\sf PROG}(\mathtt{S},\mathtt{T})$ of programs of length at most $\mathtt{S}$ that run in time at most $\mathtt{T}$ (on inputs we are considering). Typically the program length $\mathtt{S}$ will be much lower than the runtime $\mathtt{T}$. Can we learn this class with sample complexity $O(\mathtt{S})$ (or perhaps ${\mathcal{P}oly}(\mathtt{S})$, or with a mild dependence on $\mathtt{T}$ such as ${\mathcal{P}oly}(\mathtt{S}\log\mathtt{T})$), and training time ${\mathcal{P}oly}(\mathtt{T})$? Can we do this using time-invariant autoregressive learning? Can we have a (natural and simple?) base class ${\mathcal{F}}$ for every $\mathtt{S}$ and $\mathtt{T}$ such that:
@@
    We have ${\sf PROG}(\mathtt{S},\mathtt{T})\subseteq{\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}oly}(\mathtt{T})}}$;
@@
    ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}oly}(\mathtt{T})}}$ is learnable with ${\mathcal{P}oly}(\mathtt{S})$ or perhaps ${\mathcal{P}oly}(\mathtt{S}\log\mathtt{T})$ samples;
@@
    And ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}oly}(\mathtt{T})}}$ is ${\sf CoT}$-learnable in time ${\mathcal{P}oly}(\mathtt{T})$?
@@
One reason time-invariant iterated linear thresholds, $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ are not sufficient for universal expressibility, as defined above, is that they attend to only a limited context window. That is, the output $f_{{\bm{w}}}({\bm{x}})$ of a linear threshold generator $f_{{\bm{w}}}\in\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}$ depends only on the last $d$ tokens ${\bm{x}}[-d:]$ of its input ${\bm{x}}$—we refer to this as having a “context length” of $d$. When applying $f_{{\bm{w}}}$ ly for $T\gg d$ steps to obtain ${\bm{z}}=f_{{\bm{w}}}^{{\sf CoT}\textnormal{-}{T}}$, we have that ${\bm{z}}[t:]$ only depends on ${\bm{z}}[t-d:t-1]$, i.e. on a finite state of $d$ bits, and we cannot expect to capture computations requiring more than $d$ bits of memory. Since the sample complexity of learning is also linear in $d$, we see that with iterated linear thresholds the sample complexity of learning a target $h$ must be at least linear in the space complexity of computing $h$, and cannot depend only (or primarily) on the program length.
To achieve our desiderata on universal expressibility (as defined above) with sample complexity (nearly) independent of program runtime (and space), we therefore need generator base classes with a long context window (i.e. where $f({\bm{x}})$ does not depend only on a short suffix of ${\bm{x}}$) yet small complexity. One simple candidate is the class of $k$-sparse $d$-dimensional linear thresholds, which has log cardinality (and thus sample complexity) $O(k^{2}+k\log d)$ but context length $d$ (see Section D.1). However, we do not see how to achieve universal expressibility with sparse linear thresholds and suspect this is not possible. In the next section, we construct a base class that does directly allow for universal expressibility.

### v2

@@ 3.3 Computational Complexity
We now investigate the computational complexity of implementing the rules $\textsc{Cons}_{{\sf e2e}}$ and $\textsc{Cons}_{{\sf CoT}}$. As discussed in Section˜2, while all our sample complexity guarantees are for learning over the entire ${\mathcal{X}}=\Sigma^{*}$, when discussing computational complexity, the length of the input will be important. We consider a hierarchy of base classes ${\mathcal{F}}_{d}$ with some size parameter $d$. If the base class ${\mathcal{F}}_{d}$ itself is computationally hard to learn, we cannot generally hope for the iterated classes to be easier. But what can we say if ${\mathcal{F}}_{d}$ itself is computationally easy to learn? Can we then implement $\textsc{Cons}_{{\sf e2e}}$ and $\textsc{Cons}_{{\sf CoT}}$ efficiently?
@@
Consider any hypothesis class $\mathcal{F}$} over an alphabet set $\Sigma$. The rule $\textsc{Cons}_{{\sf CoT}}$ on $m$ examples of input length at most $n$ can be implemented by a single call to $\textsc{Cons}_{\mathcal{F}}$, with $\tilde{m}\leq m\cdot T$ samples of input length at most $n+T$, and O(m~)O($\tilde${m}) additional runtime.
@@
For a family ${\mathcal{F}}_{d}$, if $\textsc{Cons}_{{\mathcal{F}}_{d}}$ can be implemented in time polynomial in its input and the size parameter $d$, and $\mathrm{VCdim}({\mathcal{F}}_{d})\leq{\mathcal{P}\hskip-1.0ptoly}(d)$, then ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf CoT$} learnable in time ${\mathcal{P}\hskip-1.0ptoly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log\delta^{\scriptscriptstyle{-1}})$.
In the next section, we will see that the situation is much grimmer for $\sf e2e$}-learnability. Even for tractably learnable base classes, where $\textsc{Cons}_{\mathcal{F}}$ is computationally easy, implementing $\textsc{Cons}_{{\sf e2e}}$, or in fact $\sf e2e$}-learning using any other possible rule, could be computationally hard. This computational gap is perhaps the biggest advantage of Chain-of-Thought training. In Sections˜5 and 6, we will further see how to leverage the computational tractability of $\sf CoT$} learning exposed by Corollary˜3.7.
@@ 4 Autoregressive Linear Thresholds
From Theorems˜3.2 and 3.4, we have that $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf e2e$} and $\sf CoT$} learnable with sample complexities $m^{\sf e2e}\propto O(Td)$ and $m^{\sf CoT}\propto O(d\log T)$ respectively, with the disappointing $O(T)$ scaling for $\sf e2e$} learning. But in this case, even though $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}$ is specified in terms of real-valued parameters, the discreteness of the input domain actually allows us to bound its cardinality:
@@
Plugging in Lemma˜4.1 into Theorem˜3.1 we see that $\sf e2e$} learning is also possible with $T$-independent sample complexity. Combining the VC-dimension and cardinality-based sample complexities:
@@
While we do not know if the $d^{2}$ vs. $d\log T$ gap in the sample complexity (when $\log T\ll d$) between $\sf e2e$} and $\sf CoT$} learning is real, it is clear that there remains a statistical advantage over learning the time-dependent $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ (as in Malach, 2023), which requires $\Omega(d\cdot T)$ samples. The main advantage, however, of $\sf CoT$} learning over $\sf e2e$} learning here is computational, rather than statistical. Since the consistency problem $\textsc{Cons}_{\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}}$ amounts to Linear Programming feasibility and is thus easily solvable in polynomial time (Dantzig, 2016), Corollary˜3.7 ensures that we can $\sf CoT$} learn $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ with $O(d\log T\wedge d^{2})$ samples and implement the rule on inputs of length $n$ in time ${\mathcal{P}\hskip-1.0ptoly}(n,d,T,\varepsilon^{\scriptscriptstyle{-1}},\log(\delta^{\scriptscriptstyle{-1}}))$:
@@
The class $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf CoT$}-learnable in time ${\mathcal{P}\hskip-1.0ptoly}(d,T,n,1/\varepsilon,\log(1/\delta))$.
@@
Computational Complexity Not learnable in ${\mathcal{P}\hskip-1.0ptoly}(d,T,n)$, Learnable in ${\mathcal{P}\hskip-1.0ptoly}(d,T,n)$,
@@
Table 2: A summary of the results in Section˜4 about learnability of time-invariant iterated linear thresholds $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$. Sample complexities follow Corollary˜4.2. We do not know whether these bounds are tight, and it remains open whether the gap between $m^{{\sf e2e}}$ and $m^{{\sf CoT}}$ is real. In Corollary˜4.3 and Theorem˜4.4, we show the tractability of $\sf CoT$}-learning and hardness of $\sf e2e$}-learning of $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ respectively.
On the other hand, we show that $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ is not tractably $\sf e2e$}-learnable, i.e. when Chain-of-Thought is not provided. Not only is $\textsc{Cons}_{{\sf e2e}}$ hard to implement, but we prove that no learning rule can $\sf e2e$}-learn $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ in polynomial time (even improperly). To show this, we rely on the hardness of learning constant depth threshold circuits—a well-established hardness assumption in learning theory. See Section˜C.1 for a formal definition of the class of linear threshold circuits.
@@
There exists a constant $L>0$ and a polynomial $p(n)$ such that threshold circuits of depth $L$ and size $p(n)$ over $n$ binary inputs are not weakly PAC-learnable in time ${\mathcal{P}\hskip-1.0ptoly}(n)$, i.e. to constant accuracy and confidence parameters $\varepsilon<0.5,\delta<1$.
Assumption˜1 is implied by core cryptographic assumptions such as hardness of the problems of inverting the RSA encryption function, recognizing quadratic residues, and factoring Blum integers (see Kearns and Valiant, 1994, Theorem 6). The Decisional Diffie-Hellman (DDH) assumption implies hardness already of learning depth-4 circuits, and so also implies Assumption˜1 (Naor and Reingold, 1997, Krause and Lucks, 2001). An intersection of halfspaces is simply a depth-two linear threshold circuit, and hence hardness of learning the intersection of polynomially many halfspaces also implies Assumption˜1. This is in turn implied by hardness of the unique shortest vector problem (Klivans and Sherstov, 2009), hardness of refuting random $K$-SAT formulas (Daniely and Shalev-Shwartz, 2016), and the existence of local pseudo-random generators (Daniely and Vardi, 2021).
@@
Under Assumption˜1, there is no algorithm that $\sf e2e$}-learns ${\mathcal{F}}_{\scriptscriptstyle{d,{\rm lin}}}^{{\sf e2e}\textnormal{-}{T}}$ over $\{0,1\}^{n}$ in time ${\mathcal{P}\hskip-1.0ptoly}(d,T,n)$ (even to within constants $\varepsilon<0.5,\delta<1$).
Theorem˜4.4 follows from a reduction from circuits to time-invariant iterated linear thresholds: we show that any linear threshold circuit with depth $L$ and size $s$ can be emulated using a time-invariant autoregressive linear threshold of dimension $d=O(s^{L})$, up to a fixed poly-time feature map:
@@
In Lemma˜4.5, the insistence on time-invariance, i.e. using the same linear threshold function in every step of generation, makes the construction much trickier. If instead we allowed a different linear predictor at each step of generation, as in $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ defined in Eqs. (4)–(5), we could directly encode each node in the circuit as a step of generation, with $T=s$ generation steps and dimensionality $d=s+n$. This was the approach taken by Malach (2023) when studying time-dependent autoregressive generation. We leave it as an open question whether it is possible to improve the time-invariant construction of Lemma˜4.5, and reduce the required dimensionality, and especially the dependence on depth. In any case, Lemma˜4.5 is sufficient for obtaining the hardness result of Theorem˜4.4.999Improving the construction will only enable basing hardness on slightly weaker assumptions, such as hardness of learning log-depth circuits, or even existence of one-way functions.
@@ 5 Expressivity and Universality
One might possibly view Lemma˜4.5 as a positive result about the expressive power of autoregressive time-invariant linear thresholds, suggesting they are “universal” and can express any computable function. In this sense Lemma˜4.5 is disappointing, yielding a sample complexity of $\text{size}^{\Omega(\text{depth})}$ for learning circuits. Not only is this exponential in depth, but even for very shallow circuits, e.g. $L=2$, the sample complexity is much larger than the size of the circuit and of the sample complexity of learning the circuit directly.
@@
Should this motivate us to attempt and improve the construction in Lemma˜4.5? If our goal is to express circuits, the time-dependent class $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ is more appropriate, and we will not be able to improve over the sample complexity it provides for learning circuits (Malach, 2023, Bartlett and Maass, 2003). Circuits are inherently not “time-invariant”, and their size and number of parameters are always (at least) linear in the input length, and more importantly, in the time of the computation involved.
@@
What form of “universal” expressive power do we want then? What would we want the sample complexity to depend on? One type of desirable universality is to learn all computable functions with sample complexity proportional to their runtime. More precisely, letting $\mathsf{TIME}(\mathtt{T})$ denote the class of functions computable in time $\mathtt{T}$, we wish to learn $\mathsf{TIME}(\mathtt{T})$ with ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ samples. All such functions can be expressed as circuits of size $\tilde{O}(\mathtt{T})$ (Arora and Barak, 2009), and hence feed-forward neural nets with $E={\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ edges. Since, from a sample complexity perspective, neural nets (e.g. with threshold activations) can be learned from $\tilde{O}(E)$ samples (Anthony and Bartlett, 2009), neural nets already provide for “universal” learning in this sense: minimizing the loss on a neural net of size ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ would allow us to learn $\mathsf{TIME}(\mathtt{T})$ using ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ samples. The important caveat here, of course, is that learning a neural network is not computationally tractable, and we cannot actually minimize the loss on a neural net in polynomial time.
@@
Time-Dependent Autoregressive Chain-of-Thought learning with linear thresholds, i.e. using $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$ is also “universal” in the same way as feed-forward networks: it allows expressing the class $\mathsf{TIME}(\mathtt{T})$ with dimensionality and generation length $d,T=O(\mathtt{T})$, and thus learning with ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$ samples. The advantage here, as pointed out by Malach (2023), is that we can consider Chain-of-Thought training as essentially providing us a peek into the inner working of the computation, and thus allowing computationally tractable learning.
@@
While the above results on time-dependent autoregressive Chain-of-Thought learning are promising, we can hope for much stronger universality than just learning with number of samples proportional to computation runtime. Consider the much broader class ${\sf PROG}(\mathtt{S})$ of programs of length at most $\mathtt{S}$. Many very-long-runtime functions have a short program; runtime usually depends also on the input length while program length does not. This is also a “time-invariant” class, since the program is independent of the input and computation length. An optimistic perspective from statistical learning theory would suggest that, due to the fact that statistical complexity scales only logarithmically in the cardinality of a hypothesis class, one could, in principle, learn ${\sf PROG}(\mathtt{S})$ with only $\log{\left\lvert{{\sf PROG}(\mathtt{S})}\right\rvert}=O(\mathtt{S})$ samples; as in the case of neural networks above, however, the caveat in the previous observation is the computational intractability. We really have no hope of learning ${\sf PROG}(\mathtt{S})$ tractably–not only is finding a short program with small error uncomputable, but also functions in ${\sf PROG}(\mathtt{S})$ might require unbounded amount of time to even evaluate, let alone learn. Such functions are anyway useless as predictors, and it would be sensible to exclude them from consideration in the first place.
@@
This leads us to consider the class ${\sf PROG}(\mathtt{S},\mathtt{T})$ of programs of length at most $\mathtt{S}$ that run in time at most $\mathtt{T}$ (on inputs we are considering). Typically the program length $\mathtt{S}$ will be much lower than the runtime $\mathtt{T}$. Can we learn this class with sample complexity $O(\mathtt{S})$ (or perhaps ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{S})$, or with a mild dependence on $\mathtt{T}$ such as ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{S}\log\mathtt{T})$), and training time ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$? Can we do this using time-invariant autoregressive learning? Can we have a (natural and simple?) base class $\mathcal{F}$} for every $\mathtt{S}$ and $\mathtt{T}$ such that:
@@
    We have ${\sf PROG}(\mathtt{S},\mathtt{T})\subseteq{\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})}}$;
@@
    ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})}}$ is learnable with ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{S})$ or perhaps ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{S}\log\mathtt{T})$ samples;
@@
    And ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{{\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})}}$ is $\sf CoT$}-learnable in time ${\mathcal{P}\hskip-1.0ptoly}(\mathtt{T})$?
@@
One reason time-invariant iterated linear thresholds, $\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}^{{\sf e2e}\textnormal{-}{T}}$ are not sufficient for universal expressibility (as defined above) is that they attend to only a limited context window. That is, the output $f_{{\bm{w}}}({\bm{x}})$ of a linear threshold generator $f_{{\bm{w}}}\in\mathcal{F}_{\scriptscriptstyle d,{\rm lin}}$ depends only on the last $d$ tokens ${\bm{x}}[-d:]$ of its input $\bm{x}$}—we refer to this as having a “context length” of $d$. When applying $f_{{\bm{w}}}$ autoregressively for $T\gg d$ steps to obtain ${\bm{z}}=f_{{\bm{w}}}^{{\sf CoT}\textnormal{-}{T}}$, we have that ${\bm{z}}[t:]$ only depends on ${\bm{z}}[t-d:t-1]$, i.e. on a finite state of $d$ bits, and we cannot expect to capture computations requiring more than $d$ bits of memory. Since the sample complexity of learning is also linear in $d$, we see that with iterated linear thresholds the sample complexity of learning a target $h$ must be at least linear in the space complexity of computing $h$, and cannot depend only (or primarily) on the program length. To achieve our desiderata on universal expressibility (as defined above) with sample complexity (nearly) independent of program runtime (and space), we therefore need generator base classes with a long context window (i.e. where $f({\bm{x}})$ does not depend only on a short suffix of $\bm{x}$}) yet small complexity. [^10] One simple candidate to consider is the class of sparse linear thresholds (studied in Section D.1), which has a large context window of $d$ but its sample complexity may be only logarithmic in the context size. We see that this base class indeed satisfies two of the three desiderata, however, its expressive power remains unclear.